Show that the velocity of impact at the ground of an object thrown vertically upward is equal to the initial velocity of the upward throw.


1
Expert's answer
2021-03-24T19:44:19-0400

Let "v" is the velocity of the upward throw, "u" is the velocity of impact at the ground.

"0=v-gt," "\implies" "t=\frac vg,"

"s=vt-\frac{gt^2}{2}=\frac{v^2}{g}-\frac{v^2}{2g}=\frac{v^2}{2g};"


"s=\frac{gt'^2}{2}\implies t'=\sqrt{\frac{2s}{g}}=\sqrt{\frac{v^2}{g^2}}=\frac vg=t,"

"u=gt'=g\cdot \frac vg=v\implies u=v."